Paul is making a potion. He needs a concentration of 1.55 mil/L to put into a 500mL bottle. If he has a bottle of concentrate at 4.90 mol/L outline how he will make the dilute solution he needs
1
Expert's answer
2020-04-09T13:11:44-0400
"n_{solute} = cV = 1.55*0.5 = 0.775mol"

"V_2 = n_{solute}\/c = 0.775\/4.90 = 0.158L = 158mL"

"\Delta V = V_1 - V_2 = 500 - 158 = 342 mL"

Accordingly, it is necessary to bring a solution with a concentration of 4.90 M with a volume of 342 mL to a mark of 500 mL with water.